OpenStudy (anonymous):

Solve the polynomial equation solve x^3 + 4x^2 +x+ 6 = 0

hartnn (hartnn):

u sure the question is correct ? that equation has horrible roots!

OpenStudy (raden):

maybe for the constanta -6 not +6

OpenStudy (anonymous):

no, I checked its actually +6.

OpenStudy (sirm3d):

one root is \[\sqrt[3]{\frac{ 2\sqrt{43} }{ 3 }-\frac{ 127 }{7 }}+\sqrt[3]{-\frac{ 2\sqrt{43} }{ 3 }-\frac{ 127 }{ 7 }}-\frac{ 4 }{ 3 }\]

OpenStudy (raden):

hmmm, there are no integer solution it can solve by interpolation method, but too long so, better use wolfram like @hartnn said ^^